myininaya (myininaya):

for supercrazy: how to solve \[3^{4x-7}=4^{2x+3}\]

myininaya (myininaya):

Do you remember \[\ln(x^r)=r \ln(x)\]

myininaya (myininaya):

i got stood up

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

those loger-epo rules make more sense when defined by calculus

OpenStudy (anonymous):

@Myini-Did I do something? o.o Sorry I will delete my comment.

myininaya (myininaya):

No this person wanted to know how to do this problem in chat and I said was going to post it here and she never came. :(

OpenStudy (anonymous):

@Myini-You did your work nevertheless, shows you're a sincere person :)

OpenStudy (anonymous):

Just to practice latex \[\begin{align}3^{4x-7}=4^{2x+3} &<=> (4x-7)ln(3) &&=(2x+3)ln(4) \\ &<=> x4ln(3)-7ln(3) &&= x2ln(4)+3ln(4) \\ &<=> x(4ln(3)-2ln(4)) &&= 3ln(4)+7ln(3) \\ & <=> x &&={ 3ln(4)+7ln(3) \over 4ln(3)-2ln(4)} \\ or\ & x={ln(4^3.3^7) \over ln(3^4.4^2) }&& = {ln(139968)\over ln(1296)}\approx 1.65\end{align} \]

myininaya (myininaya):

Pretty mkone!

myininaya (myininaya):

well i think i see something just a little off on your solution

myininaya (myininaya):

\[4 \ln(3)-2 \ln(4)=\ln(3^4)-\ln(4^2)=\ln(\frac{3^4}{4^2})\]

OpenStudy (anonymous):

I'm not getting what's off is it the presentation. If so I agree \(ln({3^4 \over 4^2}) \) looks better but then what about \(2ln({3^2 \over 4}) \) ?